# DocLayLLM: An Efficient and Effective Multi-modal Extension of Large Language Models for Text-rich Document Understanding

### v2

@@ Abstract
Text-rich document understanding (TDU) refers to analyzing and comprehending documents containing substantial textual content. With the rapid evolution of large language models (LLMs), they have been widely leveraged for TDU due to their remarkable versatility and generalization. In this paper, we introduce DocLayLLM, an efficient and effective multi-modal extension of LLMs specifically designed for TDU. By integrating visual patch tokens and 2D positional tokens into LLMs and encoding the document content using the LLMs themselves, we fully take advantage of the document comprehension capability of LLMs and enhance their perception of OCR information. We have also deeply considered the role of the chain-of-thought (CoT) and innovatively proposed the techniques of CoT Pre-training and CoT Annealing. Our DocLayLLM can achieve remarkable performances with lightweight training settings, showcasing its efficiency and effectiveness. Experimental results demonstrate that our DocLayLLM surpasses existing OCR-dependent methods and also outperforms OCR-free competitors.
DocLayLLM: An Efficient and Effective Multi-modal Extension of Large Language Models for Text-rich Document Understanding
Wenhui Liao1, Jiapeng Wang1, Hongliang Li1, Chengyu Wang2, Jun Huang2, Lianwen Jin1 1South China University of Technology, Guangzhou, China 2Alibaba Group, Hangzhou, China {eelwh, eejpwang, eehongliangli}@mail.scut.edu.cn, eelwjin@scut.edu.cn {chengyu.wcy, huangjun.hj}@alibaba-inc.com
22footnotetext: Equal contributions, listed in no particular order.11footnotetext: Co-corresponding authors.
@@ 1 Introduction
Text-rich document understanding (TDU) involves analyzing and comprehending documents with extensive textual information. These documents often have complex layouts, requiring an integrated understanding of text, layout, and image data. Early research (Xu et al., 2020; Li et al., 2021a; Appalaraju et al., 2021; Li et al., 2021b; Wang et al., 2022; Hong et al., 2022; Luo et al., 2023) often used BERT-like (Devlin et al., 2019) Transformer encoders, designing various pre-training tasks for large-scale unlabeled document corpora and fine-tuning their models on specific downstream tasks like key information extraction (KIE) and visual question answering (VQA). However, these pre-training tasks often diverge significantly from downstream tasks, requiring the model to acquire substantial task-specific knowledge during fine-tuning. This may result in limited generalization capability across different tasks and datasets.
@@
With the rapid development of large language models (LLMs), their impressive versatility and generalization capabilities have captured widespread interest. This has led many researchers to explore LLMs to overcome the challenges of TDU (Zhang et al., 2023; Ye et al., 2023b; Tang et al., 2024; Hu et al., 2024a; Li et al., 2024b). Initially, efforts focused on employing multi-modal large language models (MLLMs) for OCR-free TDU, where document images are directly fed into MLLMs, generating responses to comprehension tasks. These MLLMs, which can process images and text inputs effectively, are among the most recent advancements in this domain. To date, several general MLLMs have been developed to manage diverse scenarios. Researchers often enhance these models by integrating additional document instruction-tuning data, aiming to improve their ability to comprehend documents. However, these models typically necessitate high-resolution images and an amount of training data to recognize text accurately in dense documents, which escalates the computational demands. Furthermore, to achieve superior performance, these models are usually trained using vast collections of image and instruction data, including the training sets of a specific dataset they are tested on, potentially affecting the evaluation of their ability to perform on previously unseen datasets.
@@
To overcome the constraints of OCR-free methods, scholars have pursued diverse methodologies for integrating OCR information into LLMs. Initial strategies (Wang et al., 2023; He et al., 2023) involved inputting coordinates of OCR bounding boxes as numeric text. However, this textual representation of numbers often resulted in excessively long inputs that could decrease inference speeds and potentially degrade model performance. In response, many researchers Kim et al. (2023); Fujitake (2024); Luo et al. (2024) have adopted an auxiliary document encoder (DE) like LayoutLMv3 (Huang et al., 2022) to process textual, visual, and layout features from documents and integrated into LLMs. Although these methods avoid high-resolution image inputs, the textual content of OCR documents must also be encoded alongside the layout and image features with the DE due to its structural limitations. However, the document comprehension capabilities of these DEs may not be on par with those of LLMs themselves, which might have led to significant fine-tuning efforts for aligning the feature spaces of DEs and LLMs during the whole training phase, demanding substantial computational resources.
@@
In response to the limitations inherent in OCR-dependent TDU, we introduce DocLayLLM, an efficient and effective multimodal extension of large language models for text-rich document understanding. Our method diverges from existing approaches that utilize an extra DE by integrating image patch tokens and 2D positional tokens with textual content into a natural language expression and processing them through an LLM. This innovative integration not only taps into the LLM’s robust text comprehension abilities but also substantially enhances its capability to perceive document layouts. Besides, by incorporating patch tokens, the model gains a rudimentary understanding of document structures, and supports image-related pre-training tasks.
To advance model performance, we have fully integrated the concept of the chain-of-thought (CoT) (Brown et al., 2020) into every stage of our training process. Firstly, we developed a novel CoT Pre-training strategy. CoT is engineered to preserve thematic coherence and logical progression during model reasoning, further enhancing the inference ability of LLM. It has demonstrated effectiveness in various subsequent studies (Wei et al., 2022; He et al., 2023; Luo et al., 2024). Nevertheless, the incorporation of CoT in the pre-training phase to boost reasoning capabilities remains underexplored. By employing CoT, we have improved the model’s comprehension ability while aligning these visual and layout features with LLM features.
Furthermore, we introduced a CoT Annealing technique. While previous research (Li et al., 2024a; Llama Team, 2024) has validated the effectiveness of annealing strategies that gradually increase the proportion of high-quality data towards the end of training, the impact of CoT on data quality had not been explored. To address this, we first propose an annealing strategy that considers data quality from a CoT perspective, further improving overall model performance.
@@
Thanks to the above design, utilizing only around 3M pre-training data and 300K supervised fine-tuning (SFT) data with an efficient tuning technology of LoRA (Hu et al., 2022), our DocLayLLM finishes the whole training process in 36 A100 days, while some TDU model could take up to 1200 A100 days (Tang et al., 2024), therefore resulting a highly lightweight extension of LLM. However, even with such low resources, our DocLayLLM surpasses existing OCR-dependent TDU methods. Furthermore, it also excels in specific tasks without fine-tuning on downstream tasks and outperforms state-of-the-art (SOTA) OCR-free methods under fair fine-tuning conditions, showcasing our efficiency and effectiveness. Our contributions are:
@@
  • We propose an efficient and effective multi-modal extension of LLMs, which has augmented LLM in understanding text-rich documents and significantly reduced training resources in need.
  • We introduce CoT Pre-training, remarkably enhancing the model’s data utilization efficiency and comprehension abilities.
@@
  • We re-evaluate data annealing from a novel perspective and introduce the CoT Annealing strategy, which further enhances the model’s efficiency in data utilization and boosts its ultimate performance.
@@ 2 Related Works
OCR-free MLLMs for Text-Rich Documents Understanding.  Many studies have attempted to directly answer the instruction for TDU tasks from document images, thereby obviating the need for OCR processing and resulting in an end-to-end paradigm. Such models are referred to as OCR-free models. Some early examples like Donut (Kim et al., 2022), UDOP (Tang et al., 2023), and OmniParser (Wan et al., 2024) have explored end-to-end document understanding solutions without relying on LLMs or MLLMs. With the rapid development of MLLMs, many new studies have begun to focus on enhancing the capabilities of general MLLMs in document understanding. Among these, LLaVAR (Zhang et al., 2023) is an early initiative that studied OCR-free MLLMs, utilizing GPT-4 (OpenAI, 2023) to generate image-text dialogue data for fine-tuning the LLaVA (Liu et al., 2023a) model. Inspired by LLaVAR, TextSquare (Tang et al., 2024) further uses Gemini (Google, 2023) to create 10M visual dialogue data covering more text-rich and other scenarios to augmente InternLM-Xcomposer2 (Dong et al., 2024). In addition, mPLUG-DocOwl 1.5 (Hu et al., 2024a) enhances the document understanding capability by integrating existing document comprehension data and converting it into vision-related dialogue data. Moreover, researchers have also explored in terms of image encoding and image-text alignment in MLLMs. For instance, UReader Ye et al. (2023b) and Monkey Liu et al. (2024) empower MLLMs to handle high-resolution images by segmenting these images into multiple sub-images. Then they concatenate the independently encoded features of these sub-images along with a thumbnail of the original image before feeding them into the LLM. Although these methods have achieved notable success in document understanding tasks, they typically require high-resolution image inputs to identify text within images, significantly increasing computational demands. Additionally, the absence of explicit OCR information input means that these methods still have room for improvement in tasks highly dependent on OCRed positional information like KIE. Besides, to further enhance performance, these methods often train using extensive datasets, including the training sets of the datasets used for evaluation.
@@
OCR-dependent MLLMs for Text-Rich Documents Understanding.  To address the challenges of OCR-free MLLMs, some researchers have attempted to integrate OCR information from text-rich documents into LLMs to enhance their perception of OCR information. ICL-D3IE (He et al., 2023) feeds coordinates of OCR boxes into the LLM using the format of [left, top, right, bottom], while LATIN-Prompt (Wang et al., 2023) aims to model the document layout by inserting spaces and newline characters between lines of OCR-recognized text, thus providing the LLM with a document input containing rudimentary layout information. However, these methods of inputting OCR information lead to overly lengthy inputs, subsequently slowing inference and impacting model performance. Consequently, subsequent studies have explored encoding OCR information using additional encoders. Cream (Kim et al., 2023) utilizes BLIP-2 (Li et al., 2023) to encode document OCR positions, images, text, and additional object detection information, with the output of BLIP-2 fed directly into the LLM. InstructDoc (Tanaka et al., 2024) introduces an extra feature extractor to encode document OCR information, compressing it with learnable tokens before inputting it into the LLM. Additionally, both (Fujitake, 2024) and (Luo et al., 2024) adopt LayoutLMv3 (Huang et al., 2022) to encode the OCR information of documents before inputting them into the LLM. Although these methods avoid the issue of overly lengthy text inputs, they must use the extra DE to encode all the OCR textual, positional information of the document due to the structural restriction of DE. However, the additional DE might not be comparable with LLMs themselves in terms of the document comprehension capability. This might be the reason why they usually need extensive training of both the DE and the LLM to align the features encoded by the DE with those of the LLM, which consumes significant computational resources. DocLLM (Wang et al., 2024) proposes an extension where the hidden states derived from encoded OCR positional information interact with the hidden states of OCR text features encoded by LLMs within each Transformer block via cross-attention. However, as it alters the internal structure of the LLM, it could affect its generality and generalizability, thus necessitating fine-tuning before deployment. Moreover, it does not adequately consider the role of image features in TDU. Overall, designing an efficient and effective method to provide OCR information to LLMs remains a research question worthy of deep exploration.
@@ 3 DocLayLLM
DocLayLLM is an LLM-based methodology that encodes OCRed textual, visual, and positional information directly within an LLM, eliminating the need for additional DE. The training regime of DocLayLLM includes both pre-training and supervised fine-tuning phases, which are both instruction-tuning-based processes. To further enhance our model, we propose the integration of the CoT into our pre-training tasks, denoted as CoT Pre-training. Additionally, we introduce CoT Annealing, which involves the application of annealing to the data utilized in the supervised fine-tuning phase, specifically from the perspective of CoT.
@@ 3.1 Model Architecture
The overall architecture of our DocLayLLM is illustrated in Figure 1. To process a document image, we initially employ an OCR engine. This engine extracts the OCR results, which encompass the textual content, denoted as $Text_{0:N}$, and the associated bounding box, represented by $Box_{0:N}$. Here, $N$ indicates the total number of text lines detected by the OCR, and each $Box_{i}=[x_{0},y_{0},x_{1},y_{1}]$ defines the bounding box for $Text_{i}$. The bounding box coordinates correspond to the left, top, right, and bottom boundaries of the text line. After the OCR process, the document image is segmented into patches, denoted as $Patch_{0:M}$, where $M$ represents the total number of patches. Next, the OCR results and the document patches are integrated with the fixed part of the prompt and the specific $Question$ related to the document. This amalgamation forms the final prompt input to the model, presented in a natural language expression $\mathcal{T}$: Given the document patches: {$Patch_{0:M}$} and the document text contents and locations in the form of “text, [left, top, right, bottom]: {$(Text_{i},Box_{i})_{0:N}$} {$Question$}
Subsequently, the textual components of the prompt, including the fixed expression, $Text_{0:N}$, and $Question$, are processed using the LLM’s native text embedder. This generates the embeddings for the text content as follows:
@@
where $TE$ is the native text embedder, and $TextContent$ consists of the fixed expression, $Text_{0:N}$, and $Question$.
As for the $Patch_{0:M}$ and $Box_{0:N}$, they are first embedded with an extra embedder, then projected to the LLM’s feature space via a linear projector to obtain $Emb_{V}$ and $Emb_{L}$ as:
$Emb_{V}$ $=VP(VE(Patch_{0:M})),$ (2)
$Emb_{L}$ $=LP(LE(Box_{0:N})),$ (3)
@@
Through our design, we have preserved the original structure of the LLM, ensuring its generalization capability. At the same time, we have enabled the LLM to easily accept and process visual and layout features, making it better suited for understanding text-rich documents with various layout. This approach results in a simple yet efficient and effective multi-modal extension.
Task Description
Document Description Provide a brief overview of the document.
Layout Analysis Determine the layout type (e.g., Title, Figure, Paragraph, etc) and locate specific layout elements.
Mask Language Model Restore masked words in the provided OCR text.
Table 1: The training tasks during the pre-training stage.

### v3

# DocLayLLM: An Efficient Multi-modal Extension of Large Language Models for Text-rich Document Understanding

@@ Abstract
Text-rich document understanding (TDU) requires comprehensive analysis of documents containing substantial textual content and complex layouts. While Multimodal Large Language Models (MLLMs) have achieved fast progress in this domain, existing approaches either demand significant computational resources or struggle with effective multi-modal integration. In this paper, we introduce DocLayLLM, an efficient multi-modal extension of LLMs specifically designed for TDU. By lightly integrating visual patch tokens and 2D positional tokens into LLMs’ input and encoding the document content using the LLMs themselves, we fully take advantage of the document comprehension capability of LLMs and enhance their perception of OCR information. We have also deeply considered the role of chain-of-thought (CoT) and innovatively proposed the techniques of CoT Pre-training and CoT Annealing. Our DocLayLLM can achieve remarkable performances with lightweight training settings, showcasing its efficiency and effectiveness. Experimental results demonstrate that our DocLayLLM outperforms existing OCR-dependent methods and OCR-free competitors. Code and model are available at https://github.com/whlscut/DocLayLLM.
22footnotetext: Equal contribution.11footnotetext: Co-corresponding authors.
@@ 1 Introduction
Text-rich document understanding (TDU) focuses on analyzing documents with extensive text and complex layouts, requiring a combined understanding of text, layout, and images. Early research [72, 28, 33, 55, 12, 73, 1, 37, 68, 69, 36, 52, 15, 20, 65, 77, 44, 34, 2, 35] often used BERT-like [8] Transformer-based document encoders (DE), designing various pre-training tasks for large-scale unlabeled document corpora and fine-tuning their models on specific downstream tasks like key information extraction (KIE) and visual question answering (VQA) [39, 40]. However, these pre-training tasks often diverge significantly from downstream tasks, requiring the model to acquire substantial task-specific knowledge during fine-tuning. This may result in limited generalization capability across different tasks and datasets.
@@
With the rapid development of large language models (LLMs), their impressive versatility and generalization capabilities have captured widespread interest [74]. This has led many researchers [79, 76, 61, 17, 38] to overcome the challenges of TDU with LLMs. Initially, efforts focused on employing multi-modal large language models (MLLMs) for OCR-free TDU, where document images are fed into MLLMs, directly generating responses to comprehension tasks. These MLLMs, which can process images and text inputs, are among the most recent advancements in this domain. To date, several general MLLMs have been developed to manage diverse scenarios, and researchers often build their TDU models based on these general models. However, these TDU models typically require high-resolution image inputs to recognize dense text in documents and a large volume of training data to effectively develop this capability, both of which bring significant computational demands. Furthermore, within the extensive collections of image and instruction data used to train these models, the training data for downstream evaluation tasks was often included, which might potentially impact the assessment of their generalization to previously unseen datasets.
@@
To overcome the constraints of OCR-free methods, scholars have pursued diverse methodologies for integrating OCR information into LLMs. Initial strategies [70, 14] involved inputting coordinates of OCR bounding boxes as numeric text. However, this textual representation of numbers often resulted in excessively long inputs that could decrease inference speeds and potentially degrade model performance. In response, many researchers [24, 11, 45] have adopted an auxiliary document encoder like LayoutLMv3 [20] to process textual, visual, and layout features from documents and integrated into LLMs. Although these methods avoid high-resolution image inputs, the textual content of documents must also be encoded alongside the layout and image features within the DE due to its structural limitations. However, the document comprehension capabilities of these DEs may not be on par with those of LLMs themselves, which might have led to significant fine-tuning efforts for aligning the feature spaces of DEs and LLMs to improve performances during the whole training phase, demanding substantial computational resources.
@@
To address the limitations of OCR-dependent TDU methods, we introduce DocLayLLM, an efficient multimodal extension of large language models for text-rich document understanding. Our method diverges from existing approaches that utilize an extra DE by integrating the document’s image patch tokens and 2D positional tokens with textual content into a natural language expression, and then encoding them through a unified LLM. This innovative integration not only taps into the LLM’s inherent text comprehension abilities but also allows it to perceive document layouts. Besides, by incorporating patch tokens, the model supports image-related pre-training tasks and further gains a rudimentary understanding of document structures.
To advance model performance, we have fully integrated the concept of the chain-of-thought (CoT) [3] into each stage of our training process. Firstly, we developed a novel CoT Pre-training strategy. CoT is engineered to preserve thematic coherence and logical progression during model reasoning, further enhancing the inference ability of LLM, which has proven effectiveness in various subsequent studies [71, 14, 45]. Recently, the use of Long CoT reasoning to enhance the inference capability of models has attracted significant attention [49, 63, 7], and DeepSeek-R1 [7] further emphasized the importance of utilizing CoT data for a cold start before the reinforcement learning stage. Our proposed CoT Pre-training offers a method for rapidly generating large-scale CoT data from non-QA data, thereby improving the model’s reasoning ability. By employing CoT in the data generation process and subsequently pre-training the model with the generated QA-formatted data, we have aligned the added visual and layout features with LLM’s inherent textual features while improving the model’s comprehension ability. Furthermore, we introduced a CoT Annealing technique. While previous research [31, 42] has validated the effectiveness of annealing strategies that gradually increase the proportion of high-quality data towards the end of training, the impact of CoT on data quality had not been explored. To address this, we first propose an annealing strategy that considers data quality from a CoT perspective, further improving model performance.
@@
Thanks to the above design, utilizing only around 3M pre-training data and 300K supervised fine-tuning (SFT) data with the efficient tuning technology of LoRA [18], our DocLayLLM finishes the whole training process in 36 A100 days, therefore resulting an efficient extension of LLM. Even with such low resources, our DocLayLLM surpasses existing OCR-dependent TDU methods. Furthermore, it also excels in specific tasks without fine-tuning on downstream tasks and outperforms state-of-the-art (SOTA) OCR-free methods when trained with training sets of downstream tasks as other methods, further showcasing its effectiveness. Our contributions are:
@@
  • We propose an efficient multi-modal extension of LLMs, which has augmented LLM in understanding text-rich documents and greatly reduced training resources in need.
  • We introduce CoT Pre-training, a novel approach for generating CoT data that helps enhance the comprehension capabilities of models.
@@
  • We revisit data annealing from a novel perspective and introduce the CoT Annealing strategy, which further enhances the model’s efficiency in data utilization and boosts its ultimate performance.
@@ 2 Related Works
OCR-free MLLMs for Text-Rich Documents Understanding.  To date, a number of studies have attempted to directly answer the instruction for TDU tasks from document images, thereby obviating the need for extra OCR processing and resulting in an end-to-end paradigm. Such models are referred to as OCR-free methods. Early examples like Donut [23], UDOP [62], and OmniParser [66] have explored end-to-end document understanding solutions without relying on LLMs or MLLMs. However, with the rapid development of MLLMs, many new studies have begun to focus on enhancing the capabilities of general MLLMs in document understanding. Some of them attempt to boost the document understanding ability of general MLLMs by tuning them with synthesizing TDU data generated by proprietary LLMs, including LLaVAR [79] and TextSquare [61], while mPLUG-DocOwl 1.5 [17] enhances the document understanding capability by integrating existing document comprehension data and converting it into vision-related QA-formatted data. Moreover, researchers have also explored image encoding methods to read the text more clearly. For instance, UReader [76] and Monkey [38] empower MLLMs to handle high-resolution images by splitting them into sub-images, encoding each separately, and then concatenating them. Some general MLLMs, such as LLaVA-Next [29], LLaVA-OneVision [27], and Idefics3 [26], also adopt this approach to address the challenge of TDU. Although these above methods have achieved notable success in TDU tasks, they typically require high-resolution image inputs to identify text within images, significantly increasing computational demands. Besides, to equip general MLLMs with the ability to recognize the rich text within a document, a substantial amount of training data is required, including training sets of the benchmarks tested, which resulted in a high demand for training resources and might affect the evaluation of their zero-shot inference ability.
@@
OCR-dependent MLLMs for Text-Rich Documents Understanding.  To address the challenges of OCR-free MLLMs, some researchers have attempted to integrate OCR information from text-rich documents into LLMs to avoid high-resolution image inputs. ICL-D3IE [14] feeds coordinates of OCR boxes into the LLM using a textual expression of [left, top, right, bottom], while LATIN-Prompt [70] simulates the document layout by inserting spaces and newline characters between OCR-recognized text segments, thus providing the LLM with a document input containing a rudimentary layout. Besides, LMDX [53] also input the layout in a way similar to ICL-D3IE. However, these methods of inputting OCR information in textual form lead to overly lengthy inputs, thereby slowing inference and impacting model performance. Subsequent studies have explored encoding OCR information using additional encoders. Cream [24] utilizes BLIP-2 [30] to encode document OCR positions, images, text, and additional object detection information, with the output of BLIP-2 fed into the LLM. InstructDr [60] introduces an extra feature extractor to encode document OCR information, compressing it with learnable tokens before inputting it into the LLM. Additionally,  Fujitake [11] and Luo et al. [45] adopt LayoutLMv3 [20] to encode the OCR information of documents before inputting them into the LLM. Although these methods avoid the issue of overly lengthy text inputs, they must use the extra DE to encode all the OCR textual, positional information from the document due to the structural restriction of DE. However, the additional DE might not be comparable with LLMs themselves in terms of document comprehension capability. This might be the reason why they usually need extensive training of both the DE and the LLM to align the features encoded by the DE with those of the LLM, which consumes significant computational resources. DocLLM [67] proposes an extension where the hidden states derived from OCR bounding boxes interact with those of OCR text features encoded by LLMs within each attention block via cross-attention [5, 6]. However, they use an extra projection layer at each attention block to transform and inject layout information, which might excessively interfere with the original attention computation of LLM, thereby potentially affecting LLM’s generalization capability. Moreover, it does not adequately consider the role of image features in TDU. Overall, designing an efficient method to provide OCR information to LLMs remains a research question worthy of deep exploration.
@@ 3 DocLayLLM
Figure 2: Examples of CoT Pre-training across different tasks. Red text highlights key steps in the chain of thought for each task.
DocLayLLM is an LLM-based methodology that encodes OCRed textual, visual, and positional information directly within an LLM, eliminating the need for additional DE. The training regime of DocLayLLM includes both pre-training and supervised fine-tuning phases, which are both instruction-tuning-based processes. To further enhance our model, we propose the integration of CoT into our pre-training tasks, denoted as CoT Pre-training. Additionally, we introduce CoT Annealing, which involves the application of annealing to the data utilized in the supervised fine-tuning phase, specifically from the perspective of CoT.
@@ 3.1 Model Architecture
The overall architecture of our DocLayLLM is illustrated in Figure 1. To process a document image, we first employ an OCR engine. This engine extracts the OCR results, which encompass the textual content, denoted as $Text_{1:N}$, and the associated bounding box, represented by $Box_{1:N}$. Here, $N$ indicates the total number of text segments detected by the OCR, and each $Box_{i}=[x_{0},y_{0},x_{1},y_{1}]$ defines the bounding box for $Text_{i}$. The bounding box coordinates correspond to the left, top, right, and bottom boundaries of the text segment. After the OCR process, the document image is segmented into patches, denoted as $Patch_{1:M}$, where $M$ represents the total number of patches. Next, the OCR results and the document patches are integrated with the fixed part of the prompt and the specific $Question$ related to the document. This integration forms the final prompt input to the model, presented in a natural language expression $\mathcal{T}$: Given the document patches: {$Patch_{1:M}$} and the document text contents and locations in the form of “text, [left, top, right, bottom]: {$(Text_{i},Box_{i})_{1:N}$} {$Question$}
Subsequently, the textual components of the prompt, including the fixed expression, $Text_{1:N}$, and $Question$, are processed using the LLM’s native text embedder. This generates the embeddings for the text content as follows:
@@
where $TE$ is the native text embedder of LLM, and $TextContent$ consists of the fixed expression, $Text_{1:N}$, and $Question$.
As for $Patch_{1:M}$ and $Box_{1:N}$, they are first embedded with an extra embedder, then mapped to the LLM’s feature space via a linear projector, yielding $Emb_{V}$ and $Emb_{L}$ as:
$Emb_{V}$ $=VP(VE(Patch_{1:M})),$ (2)
$Emb_{L}$ $=LP(LE(Box_{1:N})),$ (3)
@@
With our design, we have preserved the original structure of the LLM, ensuring its generalization capability. At the same time, we have enabled the LLM to easily accept and process visual and layout features, making it better suited for understanding text-rich documents with various layouts. This results in a simple yet effective multi-modal extension.